OpenStudy (anonymous):

Write the equation in slope-intercept form for the line that contains (-4,-2) and is perpendicular to the line y=1/4x-7

OpenStudy (anonymous):

m=-4 y-y1=m(x-x1) y+2=-4(x+4) y+2=-4x-16 y=-4x-18

OpenStudy (anonymous):

ok so you know that it is perpendicular to 1/4x-7=y, that means that the slope of the new equation will have to be the opposite, so in this case -4x now that you have your slope, you take your point(-4,-2) and plug it in -4(-4)=-2 16-18=-2, therefore y=-4x-18

OpenStudy (anonymous):

it must be the *opposite reciprocal*

OpenStudy (anonymous):

yup